# 1(2), Leninskie gory, Moscow 119991, Russian Federation

### v2

@@ Abstract
A method for calculating the $1/d$ expansion coefficients for solutions of integration by parts relations for Feynman integrals is presented. The idea is to use linear substitutions to transform these relations to an explicitly recursive form. The possibility of such a transformation is demonstrated for several families of massless (with one massive line) vacuum integrals up to the 7-loop level.
@@ 1 Introduction
Multiloop Feynman integrals evaluation is an important ingridient of higher-order perturbative calculations. A very efficient tool in this sence is integration by parts (IBP) relations Chetyrkin:1981qh ; Tkachov:1981wb which are used in a large number of specialized computer programs Gorishnii:1989gt ; Anastasiou:2004vj ; Smirnov:2019qkx ; Klappert:2020nbg ; Lee:2012cn ; Studerus:2009ye ; vonManteuffel:2012np ; Ruijl:2017cxj .
@@
The plan of the paper is the following. In section 2 we fix the notations and breafly recall the main points of the Baikov:1996rk ; Baikov:1996iu ; Baikov:2005nv ; Baikov:2007zza . Then in section 3 we introduce (for one-parameter case) specific linear substitutions, which we apply in section 4 to the IBP relations. In the next section 5 we consider several examples of massless (with one massive line) vacuum integral families up to the 7-loop level. Finally, in section 6 we discuss a possible extension to more complex cases.
@@ 2 Integration by parts relations for Feynman integrals
are inverse propagators. The matrix $A^{ij}_{a}$ is supposed to be symmetrical with respect to upper indices. The $1/D_{a}$ can be raised as to positive powers (“denominators” or “lines”) as non-positive (“numerators”). Summation over repeating indices is implied and in many cases (unless it’s confusing) they will be omitted. In what follows we will set $m^{2}=1$; we will also use condensed notation $(n_{1},\ldots,n_{N};d)\equiv(\underline{n})$.
@@
The matrix $\hat{A}^{ij}_{a}$ is also symmetrical with respect to $(ij)$ and is inverse to $A^{kl}_{a}$ in the sence that
@@
where we use $[{\bf I}^{-}_{a},{\bf I}^{+}_{b}]=\delta_{ab}$ to derive eq. (6).
A linear combination of $R_{ik}(\underline{n})$ with shifted $(\underline{n})$ corresponds to shifts of $(\underline{n})$ in $f(\underline{n})$ and hence to the insertion of some combination of $({\bf I}^{-}_{a},{\bf I}^{+}_{a})$ just before $f(\underline{n})$.
@@
where $f(\underline{n}_{k})$ are chosen as the most convinient for direct calculation, usually with values $(\underline{n}_{k})$ from the set $(-1,0,+1)$.
Such a reduction can be achieved as by deriving relations with an explicite recursive structure (MINCER Gorishnii:1989gt for 3-loop propagator integrals, Forcer Ruijl:2017cxj for the next loop level) as by Laporta approach Laporta:2000dsw with a number of implementations Anastasiou:2004vj ; Smirnov:2019qkx ; Klappert:2020nbg ; Lee:2012cn ; Studerus:2009ye ; vonManteuffel:2012np . In any case, reduction to irreducible integrals during actual calculations requires huge computer resources, so any progress in this field is very desirable. In this context, a direct calculation of c(n) was proposed Baikov:1996rk ; Baikov:1996iu . Indeed, if $M_{k}$ are linear independent, then $c_{k}(\underline{n})$ are rational in $d$ solutions of IBP relations and can be obtained from any other solutions set as linear combinations satisfying the condition $c_{i}(\underline{n}_{k})=\delta_{ik}$.
@@
$c_{k}(\underline{n})$$\neq$ 0only $0\quad\mbox{only if}\,\,n_{a}>0\,\,\mbox{for all}\,\,a\in H_{k},$
where $H_{k}=\{a_{1},\dots,a_{k}\}$ is the set of line numbers of the irreducible integral $M_{k}$. We will use also (for auxiliary constructions) an alternative boundary condition for $a\notin H_{k}$:
$f_{k}(\underline{n})$$\neq$ 0only $0\quad\mbox{only if}\,\,n_{a}<1\,\,\mbox{for all}\,\,a\notin H_{k}.$ (8)
@@ 3 Linear substitutions in recurrence relations: one-parameter case
where $S({\bf I})$ is any formal seria in ${\bf I}$.
We are going to study the transformation of eq. (11) after linear redefinition of $f(n)$. The idea is to replace $f(n)$ with some combination of $g(n)$ to obtain a simpler recurrence relation for $g(n)$, solve them and re-express the results in terms of $f(n)$.
@@
The recurrence relation for $g(n)$ follows from the linear combination of eqs. (11) with shifted $n$:
00 0 $=$ $e^{\lambda{\bf I}_{R}^{-}}R(n)=R({\bf I}_{f}^{-},{\bf I}_{f}^{+})e^{\lambda{\bf I}_{f}^{-}}f(n)=R({\bf I}_{f}^{-},{\bf I}_{f}^{+})g(n)$
@@
that is the infinite sum of $g(n<1)$. However, as we will see in section 4 one can choose $\lambda$ so that the $k$-term in the sum (20) has leading asymptotics $d^{-k/2}$ (see eq. (21)), therefore, by keeping enough terms in this sum and assuming that $g(n<1)$ satisfies eq. (18) we can satisfy (11) in any desired order in $1/d$. As a result, $f(n)$ can be parameterized using $g(k<1)$ for any value of $n$. In what follows, by the expression (17) in the case (16) we will mean the extension (20) for $n>0$ (in the $\lambda=0$ case when (20) is not defined we will set $f(n)=g(n)$).
@@ 4 Linear substitutions in recurrence relations: Feynman integrals
Let us now return to the general case and construct, if possible, a solution to eq. (5) for some set of propagator indices $H_{k}=\{a_{1},\dots,a_{k}\}$. We will call the index $a\in H_{k}$ “hard”. (substitutions (15) will be applied), and $a\notin H_{k}$ “soft’ (substitutions (1720)).
For any indexed object $x_{a}$ we define projections onto the “hard” and “soft” subspaces: ${\overline{x}}_{a}\equiv x_{\overline{a}}=$ (if $a\in H_{k}$ then $x_{a}$ else 00) and ${\tilde{x}}_{a}\equiv x_{\tilde{a}}=x_{a}-{\overline{x}}_{a}$; as the result $x_{a}=\overline{x}_{a}+\tilde{x}_{a}$. For $x_{a}={\bf I}_{a}^{\pm}$ we have ${\bf I}_{a}^{\pm}={\tilde{\bf I}}_{a}^{\pm}+{\overline{\bf I}}_{a}^{\pm},\,\big{[}{\tilde{\bf I}}_{a}^{\pm},{\overline{\bf I}}_{b}^{\pm}\big{]}=0.$
@@
$d/2$ $=$ $1/\alpha^{2},\quad\alpha\rightarrow 0$
@@
$f(\underline{n})$ $=$ $\exp(\alpha^{-2}\overline{\lambda}\,\overline{{\bf I}}_{g}^{-}-\tilde{\lambda}\tilde{{\bf I}}_{g}^{+})f^{\prime}(\underline{n}),$
$f^{\prime}(\underline{n})$ $=$ $\alpha^{-\sum\underline{n}}g(\underline{n})$ (21)
will change $\overline{{\bf I}}_{f,a}^{+}=\alpha^{-2}(\alpha\overline{{\bf I}}_{g,a}^{+}+\overline{\lambda}_{a}),\,\tilde{{\bf I}}_{f,a}^{-}=\alpha\tilde{{\bf I}}_{g,a}^{-}+\tilde{\lambda}_{a}$ and therefore eq. (5) will be transformed into
00 0 $=$ $\Big{(}\delta_{ik}-\hat{A}^{kl}(\alpha\,{\bf I}^{-}_{g}+\tilde{\lambda}+\mu)A^{li}\big{(}{\alpha\,\bf I}_{g}^{+}+\overline{\lambda}\big{)}\Big{)}\,g(\underline{n})\,$ (22)
@@
Compared to eq. (5), we have changed the order of the indices of the matrix $A^{kl}_{a}$ (since it is symmetric). Eqs. (22) generate relations for the expansion coefficients $g_{j}(\underline{n})$ in $g(\underline{n})=\sum_{j=0}^{\infty}\alpha^{j}\ g_{j}(\underline{n})$ (in what follows we omit the subscript $g$ in ${\bf I}^{\pm}_{g}$):
00 0 $=$ $\big{(}\delta_{ik}-\hat{A}^{kl}(\tilde{\lambda}+\mu)A^{li}\overline{\lambda}\big{)}g_{j}(\underline{n})$ (23)
@@
Equations for $g_{0}(\underline{n})$ do not contain ${\bf I}_{a}^{\pm}$ operators and in general leads to $g_{0}(\underline{n})=0$ (and as a result $g_{j}(\underline{n})=0$), therefore, to obtain non-trivial solutions it is necessary to choose $\lambda_{a}$ satisfies the equation
@@
Eq. (23) with the condition (24) becomes simpler:
00 0 $=$ $\big{(}\hat{A}^{kl}(\tilde{\lambda}+\mu)A^{li}{\bf I}^{+}+\hat{A}^{kl}{\bf I}^{-}A^{li}\overline{\lambda}\big{)}g_{j}(\underline{n})$ (25)
@@
We are going to use eq. (25) to reduce $g_{j}(\underline{n})$ (up to $g_{j-1}(\underline{n})$ terms) to “corner” values: “hard” $\overline{n}_{a}\to 1$ and “soft” $\tilde{n}_{a}\to 0$ (assuming “soft” $\tilde{\lambda}_{a}\neq 0$, see the end of section 3). This would be possible if the $g_{j}(\underline{n})$ part of eq. (25) could be solved explicitly with respect to the operators ($\overline{\bf I}_{a}^{+},\tilde{\bf I}_{a}^{-}$), increasing the distance from the corner. Let us multiply eqs. (25) by $(\hat{A}_{a}A\overline{\lambda})^{ik}$:
00 0 $=$ $\Big{(}\mbox{Tr}\big{(}\hat{A}_{a}A\overline{\lambda}\,\hat{A}(\tilde{\lambda}+\mu)A{\bf I}^{+}\big{)}+\mbox{Tr}\big{(}\hat{A}_{a}A\overline{\lambda}\,\hat{A}{\bf I}^{-}A\overline{\lambda}\big{)}\Big{)}g_{j}(\underline{n})+\mbox{Tr}(\dots)g_{j-1}(\underline{n})$ (26)
@@
Here $\mbox{Tr}(\cdot)$ means a contraction of outer indices $(ik)$; we also use the condition (24) and $\mbox{Tr}(\hat{A}_{a}A_{b})=\delta_{ab}$. In the case of “soft” $\tilde{a}$, the $g_{j}(\underline{n})$ part of eq. (26) depends on increasing $\tilde{\bf I}_{a}^{-}$ and decreasing $(\tilde{\bf I}_{b}^{+},\,\overline{\bf I}_{b}^{-})$ operators and if
@@
$\tilde{\bf I}_{a}^{-}$$\rightarrow$ $(\tilde{\bf I}_{b}^{+},\,\overline{\bf I}_{b}^{-}).$ (28)
For “hard” $\overline{a}$ eq. (26) explicitly expresses the remaining increasing operators $\overline{\bf I}_{a}^{+}$ in terms of others $(\tilde{\bf I}_{b}^{-},\,\tilde{\bf I}_{b}^{+},\,\overline{\bf I}_{b}^{-})$ and due to eq. (28) in terms of decreasing $(\tilde{\bf I}_{b}^{+},\,\overline{\bf I}_{b}^{-})$ only, so as the result we have an explicite reduction to “corner” values
$(\tilde{\bf I}_{a}^{-},\,\overline{\bf I}_{a}^{+})$$\rightarrow$ $(\tilde{\bf I}_{b}^{+},\,\overline{\bf I}_{b}^{-}).$ (29)
The $L(L+1)/2$ equations (29) are linear combinations of the $L^{2}$ eqs. (26) (and hence of the $L^{2}$ eqs. (25)), so we can expect additional linear relations on ${\bf I}_{a}^{\pm}$, which can lead to a contradiction. Fortunately if (27) and hence (29) are valid this is not the case. Indeed, eqs. (26) are the combinations of eqs. (29). Then the $g_{j}(\underline{n})$ parts of eqs. (25) are the linear combinations of the $g_{j}(\underline{n})$ parts of eqs. (26) with the coefficients $(\hat{A}(\tilde{\lambda}+\mu)A_{a})^{ki}$, so substituting (29) into eq. (25) nullifies this part and no additional relations appear.
@@
Eqs. (24) leads to several useful consequences. First, $\mbox{Tr}(\cdot)$ of eqs. (24) due to eq. (4) leads to
@@
11 11 $=$ $\det\big{(}\hat{A}(\tilde{\lambda}+\mu)\big{)}\det(A\overline{\lambda})$ (31)
@@
where $A^{\mbox{\small Adj}}\equiv\det(A)\,A^{-1}$, $\partial_{a}\det(A\lambda)\equiv\partial_{x_{a}}\det(A^{ik}_{a}x_{a})|_{x_{a}=\lambda_{a}}$. For “soft” $\tilde{a}$ eq. (33) gives
$\partial_{\tilde{a}}\det(\hat{A}(\tilde{\lambda}+\mu))$ $=$ $0$ (34)
Using eq. (32) we can reexpress eq. (27) via $\tilde{\lambda}_{a}$:
00 0 $\neq$ $\det_{\tilde{a}\tilde{b}}\mbox{Tr}\big{(}\hat{A}_{\tilde{a}}A\overline{\lambda}\,\hat{A}_{\tilde{b}}A\overline{\lambda}\big{)}$ (35)
@@
$\propto$ $\det_{\tilde{a}\tilde{b}}\Big{(}\partial_{\tilde{a}}\partial_{\tilde{b}}\det\big{(}\hat{A}(\tilde{\lambda}+\mu)\big{)}\Big{)}/\det\big{(}\hat{A}(\tilde{\lambda}+\mu)\big{)},$
where in the third line we use eq. (34) and standard matrix formulas

### v3

@@ Abstract
A method for calculating the $1/d$ expansion coefficients for solutions of integration by parts relations for Feynman integrals is presented. The idea is to use linear substitutions to transform these relations to an explicitly recursive form. A possible type of such substitutions is proposed for the case of vacuum integrals. Its applicability is shown for several families of massless (with one massive line) vacuum integrals up to the 7-loop level.
@@ 1 Introduction
Multiloop Feynman integrals evaluation is an important ingredient of higher-order perturbative calculations. A very efficient tool in this sense is integration by parts (IBP) relations Chetyrkin:1981qh ; Tkachov:1981wb which are used in a large number of specialized computer programs Gorishnii:1989gt ; Anastasiou:2004vj ; Smirnov:2019qkx ; Klappert:2020nbg ; Lee:2012cn ; Studerus:2009ye ; vonManteuffel:2012np ; Ruijl:2017cxj .
@@
The plan of the paper is the following. In section 2 we fix the notations and briefly recall the main points of the Baikov:1996rk ; Baikov:1996iu ; Baikov:2005nv ; Baikov:2007zza . Then in section 3 we introduce (for one-parameter case) specific linear substitutions, which we apply in section 4 to the IBP relations. In the next section 5 we consider several examples of massless (with one massive line) vacuum integral families up to the 7-loop level. Finally, in section 6 we discuss a possible extension to more complex cases.
@@ 2 Integration by parts relations for Feynman integrals
are inverse propagators. The matrix $A^{ij}_{a}$ is supposed to be symmetrical with respect to upper indices. The $1/D_{a}$ can be raised as to positive powers (“denominators” or “lines”) as non-positive (“numerators”). Summation over repeating indices is implied and in many cases (unless it’s confusing) it will be omitted. In what follows we will set $m^{2}=1$; we will also omit $d$ in the arguments and use the shorthand $(n_{1},\ldots,n_{N})\equiv(\underline{n})$.
@@
The matrix $\hat{A}^{ij}_{a}$ is also symmetrical with respect to $(ij)$ and is inverse to $A^{kl}_{a}$ in the sense that
@@
where we use ${\bf I}^{-}_{a},{\bf I}^{+}_{b}]=\delta_{ab}$ to derive eq. (6).
[…]
@@
where $f(\underline{n}_{k})$ are chosen as the most convenient for direct calculation, usually with values $(\underline{n}_{k})$ from the set (1,0,+1)(-1,0,+1).
Such a reduction can be achieved as by deriving relations with an explicitly recursive structure (MINCER Gorishnii:1989gt for 3-loop propagator integrals, Forcer Ruijl:2017cxj for the next loop level) as by Laporta approach Laporta:2000dsw with a number of implementations Anastasiou:2004vj ; Smirnov:2019qkx ; Klappert:2020nbg ; Studerus:2009ye ; vonManteuffel:2012np . In any case, reduction to irreducible integrals during actual calculations requires huge computer resources, so any progress in this field is very desirable. In this context, a direct calculation of c(n) was proposed Baikov:1996rk ; Baikov:1996iu . Indeed, if $M_{k}$ are linearly independent, then $c_{k}(\underline{n})$ are rational in $d$ solutions of IBP relations and can be obtained from any other solutions set as linear combinations satisfying the condition $c_{i}(\underline{n}_{k})=\delta_{ik}$.
@@
$c_{k}(\underline{n})$ $\neq$ $0\quad\mbox{only if}\,\,n_{a}>0\,\,\mbox{for all}\,\,a\in H_{k},$
where $H_{k}=\{a_{1},\dots,a_{N_{k}}\}$ is the set of line numbers of the irreducible integral $M_{k}$ and $N_{k}$ is the total number of lines in $M_{k}$. In some cases we will also use an alternative boundary condition for $a\notin H_{k}$:
$f(\underline{n})$ $\neq$ $0\quad\mbox{only if}\,\,n_{a}<1\,\,\mbox{for all}\,\,a\notin H_{k}.$ (8)
@@ 3 Linear substitutions in recurrence relations: one-parameter case
where $S({\bf I})$ is any formal series in $\bf I$}.
We are going to study the transformation of eq. (11) after linear redefinition of $f(n)$. The idea is to replace $f(n)$ with some combination of $g(n)$ to obtain a simpler recurrence relation for $g(n)$, solve it and re-express the results in terms of $f(n)$.
@@
The recurrence relation for $g(n)$ follows from the linear combination of eqs. (11) with shifted $n$ (${\bf I}_{R}^{-}R(n)\equiv R(n-1)$):
$0$ $=$ $e^{\lambda{\bf I}_{R}^{-}}R(n)=R({\bf I}_{f}^{-},{\bf I}_{f}^{+})e^{\lambda{\bf I}_{f}^{-}}f(n)=R({\bf I}_{f}^{-},{\bf I}_{f}^{+})g(n)$
@@
that is the infinite sum of $g(n<1)$. However, as we will see in section 4 one can choose $\lambda$ so that the $k$-term in the sum (20) has leading asymptotics $d^{-k/2}$ (see eq. (21)), therefore, by keeping enough terms in this sum and assuming that $g(n<1)$ satisfies eq. (18) we can satisfy (11) in any desired order in $1/d$. As a result, $f(n)$ can be parameterized using $g(k<1)$ for any value of nn. [^1] Note that a direct extension of eq. (17) for $n>0$ produces an infinite sum over $k$ of terms with undesirable leading asymptotic $d^{k/2}$ (see eq. (21)). In what follows, by the expression (17) in the case (16) we will mean the extension (20) for $n>0$ (in the $\lambda=0$ case when (20) is not defined we will set $f(n)=g(n)$).
@@ 4 Linear substitutions in recurrence relations: Feynman integrals
Let us now return to the general case and construct, if possible, a solution to eq. (5) for some set of propagator indices $H_{k}=\{a_{1},\dots,a_{N_{k}}\}$. We will call the index $a\in H_{k}$ “hard”. (substitutions (15) will be applied), and $a\notin H_{k}$ “soft’ (substitutions (1720)).
For any indexed object $x_{a}$ we define projections onto the “hard” and “soft” subspaces: ${\overline{x}}_{a}\equiv x_{\overline{a}}=$ (if $a\in H_{k}$ then $x_{a}$ else 0) and ${\tilde{x}}_{a}\equiv x_{\tilde{a}}=x_{a}-{\overline{x}}_{a}$; as the result $x_{a}=\overline{x}_{a}+\tilde{x}_{a}$. For $x_{a}={\bf I}_{a}^{\pm}$ we have ${\bf I}_{a}^{\pm}={\tilde{\bf I}}_{a}^{\pm}+{\overline{\bf I}}_{a}^{\pm},\,\big[{\tilde{\bf I}}_{a}^{\pm},{\overline{\bf I}}_{b}^{\pm}\big]=0.$
@@
$d/2$ $=$ $1/\alpha^{2},\quad\alpha\rightarrow 0.$
@@
$f(\underline{n})$ $=$ $\exp(\alpha^{-2}\overline{\lambda}_{a}\,\overline{{\bf I}}_{f^{\prime},a}^{-}-\tilde{\lambda}_{a}\tilde{\bf I}_{f^{\prime},a}^{+})f^{\prime}(\underline{n}),$
$f^{\prime}(\underline{n})$ $=$ $\alpha^{N_{k}-\sum_{a}n_{a}}g(\underline{n})$ (21)
will change
$\overline{{\bf I}}_{f,a}^{+}=\alpha^{-2}(\alpha\overline{{\bf I}}_{g,a}^{+}+\overline{\lambda}_{a}),\quad\tilde{{\bf I}}_{f,a}^{-}=\alpha\tilde{{\bf I}}_{g,a}^{-}+\tilde{\lambda}_{a},\quad\tilde{I}_{f}^{+}=\alpha^{-1}\tilde{I}_{g}^{+},\quad\overline{I}_{f}^{-}=\alpha\overline{I}_{g}^{-},$
and therefore eq. (5) (multiplied by $\alpha^{2}$) will be transformed into
$0$ $=$ $\Big(\delta_{ik}-\hat{A}^{kl}(\alpha\,{\bf I}^{-}_{g}+\tilde{\lambda}+\mu)A^{li}\big({\alpha\,\bf I}_{g}^{+}+\overline{\lambda}\big)\Big)\,g(\underline{n})\,$ (22)
@@
Here and in some places below we omit the “propagator” indices over which the summation is performed, assuming that matrices $A,\hat{A}$ without a subscript are summed with the object located on the right. Also, compared to eq. (5), we have changed the order of the upper indices of the matrix $A^{kl}_{a}$ (since it is symmetrical).
Eqs. (22) generate relations for the expansion coefficients $g_{j}(\underline{n})$ in $g(\underline{n})=\sum_{j=0}^{\infty}\alpha^{j}\ g_{j}(\underline{n})$:
$0$ $=$ $\big(\delta_{ik}-\hat{A}^{kl}(\tilde{\lambda}+\mu)A^{li}\overline{\lambda}\big)g_{j}(\underline{n})$ (23)
@@
From here on we omit the indices $f,f^{\prime},g$ in $\bf I$$^{\pm}$ if the “default” action is implied.
Equations for $g_{0}(\underline{n})$ do not contain ${\bf I}_{a}^{\pm}$ operators and in general lead to $g_{0}(\underline{n})=0$ (and as a result $g_{j}(\underline{n})=0$), therefore, to obtain non-trivial solutions it is necessary to choose $\lambda_{a}$ satisfies the equations
@@
Eqs. (23) with the condition (24) become simpler:
$0$ $=$ $\big(\hat{A}^{kl}(\tilde{\lambda}+\mu)A^{li}{\bf I}^{+}+\hat{A}^{kl}{\bf I}^{-}A^{li}\overline{\lambda}\big)g_{j}(\underline{n})$ (25)
@@
We are going to use eqs. (25) to reduce $g_{j}(\underline{n})$ (up to $g_{j-1}(\underline{n})$ terms) to “corner” values: “hard” $\overline{n}_{a}\to 1$ and “soft” $\tilde{n}_{a}\to 0$ (assuming “soft” $\tilde{\lambda}_{a}\neq 0$, see the end of section 3). This would be possible if the $g_{j}(\underline{n})$ part of eqs. (25) could be solved explicitly with respect to the operators ($\overline{\bf I}_{a}^{+},\tilde{\bf I}_{a}^{-}$), increasing the distance from the corner. Let us multiply eqs. (25) by $(\hat{A}_{a}A\overline{\lambda})^{ik}$:
$0$ $=$ $\Big(\mbox{Tr}\big(\hat{A}_{a}A\overline{\lambda}\,\hat{A}(\tilde{\lambda}+\mu)A{\bf I}^{+}\big)+\mbox{Tr}\big(\hat{A}_{a}A\overline{\lambda}\,\hat{A}{\bf I}^{-}A\overline{\lambda}\big)\Big)g_{j}(\underline{n})+\mbox{Tr}(\dots)g_{j-1}(\underline{n})$ (26)
@@
Here we omit the repeated “loop” indices, $\mbox{Tr}(\cdot)$ means the contraction of outer indices $(ik)$; we also use the condition (24) and $\mbox{Tr}(\hat{A}_{a}A_{b})=\delta_{ab}$. In the case of “soft” $\tilde{a}$, the $g_{j}(\underline{n})$ part of eqs. (26) depends on increasing $\tilde{\bf I}_{a}^{-}$ and decreasing $(\tilde{\bf I}_{a}^{+},\,\overline{\bf I}_{a}^{-})$ operators and if
@@
$\tilde{\bf I}_{a}^{-}$ $\rightarrow$ $(\tilde{\bf I}_{a}^{+},\,\overline{\bf I}_{a}^{-}).$ (28)
For “hard” $\overline{a}$ eqs. (26) explicitly express the remaining increasing operators $\overline{\bf I}_{a}^{+}$ in terms of $(\tilde{\bf I}_{a}^{-},\,\overline{\bf I}_{a}^{-})$ and due to eqs. (28) in terms of decreasing $(\tilde{\bf I}_{a}^{+},\,\overline{\bf I}_{a}^{-})$ only, so as the result we have an explicit reduction to “corner” values
$(\tilde{\bf I}_{a}^{-},\,\overline{\bf I}_{a}^{+})$ $\rightarrow$ $(\tilde{\bf I}_{a}^{+},\,\overline{\bf I}_{a}^{-}).$ (29)
The $L(L+1)/2$ equations (29) are linear combinations (by means of eqs. (26)) of the $L^{2}$ eqs. (25), so we can expect additional linear relations on ${\bf I}_{a}^{\pm}$, which can lead to a contradiction. Fortunately if (27) and hence (29) are valid this is not the case. Indeed, eqs. (26) are the combinations of eqs. (29). Then the $g_{j}(\underline{n})$ parts of eqs. (25) are the linear combinations of the $g_{j}(\underline{n})$ parts of eqs. (26) with the coefficients $(\hat{A}(\tilde{\lambda}+\mu)A_{a})^{ki}$, so substituting (29) into eqs. (25) nullifies this part and no additional relations appear.
@@
Eqs. (24) lead to several useful consequences. First, $\mbox{Tr}(\cdot)$ of eqs. (24) due to eq. (4) leads to
@@
$1$ $=$ $\det\big(\hat{A}(\tilde{\lambda}+\mu)\big)\det(A\overline{\lambda})$ (31)
@@
where $A^{\mbox{\small Adj}}\equiv\det(A)\,A^{-1}$, $\partial_{a}\det(A\lambda)\equiv\partial_{x_{a}}\det(A^{ik}_{a}x_{a})|_{x_{a}=\lambda_{a}}$. For “soft” $\tilde{a}$ eqs. (33) give
$\partial_{\tilde{a}}\det(\hat{A}(\tilde{\lambda}+\mu))$ $=$ $0.$ (34)
Using eqs. (32) we can re-express eq. (27) via $\tilde{\lambda}_{a}$:
$0$ $\neq$ $\det_{\tilde{a}\tilde{b}}\mbox{Tr}\big(\hat{A}_{\tilde{a}}A\overline{\lambda}\,\hat{A}_{\tilde{b}}A\overline{\lambda}\big)$ (35)
@@
$\propto$ $\det_{\tilde{a}\tilde{b}}\Big(\partial_{\tilde{a}}\partial_{\tilde{b}}\det\big(\hat{A}(\tilde{\lambda}+\mu)\big)/\det\big(\hat{A}(\tilde{\lambda}+\mu)\big)\Big),$
where in the third line we use eqs. (34) and standard matrix formulas